you are considering a stock investment in one of two firms( lots of debt inc and lots of equity inc.) both of which operate in the same industry. lots of debt, inc finances its 35.25 million in assets with 35.75 million in debt and 2.50 million in equity. lots of equity, inc. finances its 35.25 million in assets with 2.50 million in debt and 32.75.
calculate the debt ratio.
calculate the equity mulitplier.
calculate the debt to equity.
Answer #1

LOTS OF DEBT INC.

1. Debt ratio = total debt/total assets

= 32.75/35.25

= 0.9291 or 92.91%

2. Equity multiplier = total assets/total shareholders equity

= 35.25/2.5

= 14.1 times

3. Debt to equity ratio = total debt/total shareholders equity

= 32.75/2.5

= 13.1 times

LOTS OF EQUITY INC.

1. Debt ratio = total debt/total assets

= 2.5/35.25

= 0.0709 or 7.09%

2. Equity multiplier = total assets/total shareholders equity

= 35.25/32.75

= 1.0763 times

3. Debt to equity ratio = total debt/total shareholders equity

= 2.5/32.75

= 0.0763 times